If a corporation’s taxable income is

It pays that amount on the base of the bracket

Plus this percentage on the excess over the base

Average tax rate at top of bracket

Up to 50,000

0

15%

15.0%

50,000-75,000

7,500

25

18.3

75,000-100,000

13,750

34

22.3

100,000-335,000

22,250

39

34.0

335,000-10,000,000

113,900

34

34.0

10,000,000-15,000,000

3,400,000

35

34.3

15,000,000-18,333,333

5,150,000

38

35.0

Over 18,333,333

6,416,667

35

35.0

1. ABC Inc. 2018 sales are $1,100,000. Operating costs (excluding depreciation) are 70% of sales. Net fixed assets are $205,000. Depreciation amounted to 15% of net fixed assets. Interest expenses are $100,000. The tax bill must be calculated using the corporate income tax table in the text, and ABC Inc. paid 8% of net income in dividends.

Prepare ABC Inc.’s income statement for 2018:

Sales 1,100,000

Operating costs (excluding depreciation) 1,100.000 X .70=770,000

EBITDA 1,100.00-770,00=330,00

Depreciation 205,000 x .15= 30,750

EBIT 330,000-30,750=299,250

Interest Expense 100,000

EBT 299,250-100,000=199,250

The answers below are what I need help with

Taxes***

Net Income

Common Dividends

Addition to Retained Earnings

Operating Cash Flow

***For Tax Rate use the corporate tax table in chapter 2 to determine the company’s tax bill, average and marginal tax Use the average tax rate for this portion of the income statement.

Average tax ________

Marginal tax ________
Answer #1

Sales 1,100,000
Operating costs (excluding depreciation) 1,100.000 X .70=770,000
EBITDA 1,100.00-770,00=330,00
Depreciation 205,000 x .15= 30,750
EBIT 330,000-30,750=299,250
Interest Expense 100,000
EBT 299,250-100,000=199250
taxable income is 199250 and it falls in range 100000 and 335000
Taxes as per table given is
=22250+((199250-100000)*39%)=60957.50
Net income=199250-60957.50=138292.50
Common dividend-8%*138292.5=11063.40
Addition to retained earnings=138292.50-11063.40=127229.10

Operating cash flow= net income+ depreciation
=138292.50+30750=169042.50

Average tax=total tax/taxable income
=60957.50/199250=30.59%
Marginal tax means tax rate for additional $1 increase in taxable income and here it is 39%